Zachary L. (SSN 123-45-6789) and Cici K. Mao (SSN 987-65-4321) reside at 520 Chestnut Street, Philadelphia, Pennsylvania 19106. They are both under 65 years old and have one dependent (their son, Oliver, age 4, SSN 111-22-3333). They do not wish to take advantage of the presidential election campaign check-off. Other relevant information includes:

Zachary’s salary as a manager $40,000

Cici’s salary as a teacher 45,000

Interest (First National Bank) 200

Withholding 5,600

Complete their Form 1040 for 2019, assuming that they are married and file jointly.
Answer #1

Answer:-

For Zachary.

Salary Taxable $ 45000

Interest (Bank) $ 200

Total taxable Salary $45200

Tax bracket 22℅

Tax owed $ 4543 from $ 0- $ 39475

Above 39474- 45200 = 5724 *22℅ = $1259

Total tax = $1259+$ 4543 = $ 5802

Tax withold $ 5600

Tax payable $ 202

For Cici

Salary $ 40000

Interest Bank $200

Total income $40200

Tax $ 4543 + (40200-39475) *22% +159

Total tax liability $4702

Tax withhold $5600

Tax refundable $ 898